A closed container in lab contains a sample of oxygen gas. When the temperature in lab is 14.3 degrees Celcius, the pressure in the container is 10.3 atm. What would the temperature in lab be (in Celcius) if the pressure in the container was 24.0 atm? Assume the mol and volume of the container are constant. Record your answer as a number with the correct significant figures.


1
Expert's answer
2021-12-17T14:37:02-0500

Temperature = "\frac{(14.3+273)×24}{10.3}=669.43-273=396.43°c"